# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ 3.2.4 Analysis of the limiting 1-dimensional process
Now that we have shown the giant component process near the fixed point $\theta_{*}n$ is well-approximated by the 1-dimensional process $Z_{t}$ of (40), we show here that this simplified process’s exit probabilities to the right and left are monotone, and oscillate on the $\sqrt{n}$ scale. In what follows, let $\tau_{\gamma}^{Z}$ be the hitting time of $\gamma\sqrt{n}$ for $Z_{t}$ from (40) and let $\tau_{-\gamma}^{Z}$ be the hitting time of $-\gamma\sqrt{n}$.
@@
The Markov chain of (40) is monotone. In particular, for every $p\in(0,1)$, there is a unique $c_{*}\in\mathbb{R}$ such that if $p_{0}=\lambda_{*}+c_{*}n^{-1/2}+o(n^{-1/2})$ and $Z_{0}\sim\mathcal{N}((p_{0}-\lambda_{*})\alpha^{\prime}(\lambda_{*})n,\sigma^{2}(\lambda_{*})n)$ then
@@
For the monotonicity, suppose $Z_{t}\leq Z_{t}^{\prime}$ and consider $Z_{t+1}$ and $Z_{t+1}^{\prime}$ generated per (40) using the same Bernoulli random variable $\varepsilon_{t+1}$ and the same pair of normal random variables. If $\varepsilon_{t+1}=0$ then $Z_{t+1}=Z_{t}\leq Z_{t}^{\prime}=Z_{t+1}^{\prime}$ and the monotonicity is preserved. If $\varepsilon_{t+1}=1$, then under this coupling
@@
This will be positive because $Z_{t}^{\prime}-Z_{t}\geq 0$ and $f^{\prime}(\theta_{*})>0$.
To see the other consequences, notice first that the process $(\bar{Z}_{t})_{t\geq 0}=(Z_{t}n^{-1/2})_{t\geq 0}$ is $n$-independent (as all of $h_{1},h_{2},\sigma_{t},f^{\prime},\beta_{\mathsf{u}}$ are $n$-independent). This implies that $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ are $n$-independent from $n$-independent initializations, which will be the case if $p_{0}=\lambda_{*}+c_{*}n^{-1/2}$. As $Z_{0}\to\gamma\sqrt{n}$, the probability $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ is easily checked to go to $1$, and as $Z_{0}\to-\gamma\sqrt{n}$, it goes to zero. Finally, to see that $c_{*}$ is $\gamma$-independent with the $o_{\gamma}(1)$, similarly observe that
@@
Suppose $X_{0}\sim G(n,p_{0}/n)$ for $p_{0}=\lambda_{*}+O(n^{-1/2})$ and let $(Z_{t})_{t}$ be the process of Definition 3.18 initialized from $Z_{0}\sim\mathcal{N}((p_{0}-\lambda_{*})\alpha^{\prime}(\lambda_{*})n,\sigma^{2}(\lambda_{*})n)$. Then the hitting probabilities $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ and $\mathbb{P}(\tau_{\gamma}^{X}<\tau_{-\gamma}^{X})$ are within $1-o_{\gamma,n}(1)$ of one another.
@@
The same holds for hitting $-\gamma\sqrt{n}$, as well as for the converse implications, i.e., that if $X$ has hit one side or the other then so has $Z$ or will in the next step. Putting these together, we deduce that for every $\varepsilon$, there exists $\gamma$ sufficiently large such that $|\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})-\mathbb{P}(\tau_{\gamma}^{X}<\tau_{-\gamma}^{X})|\leq\varepsilon+o(1)$. ∎
3.3 Proof of Theorem 1.1
Off-criticality. We begin by concluding the bounds of mixing when initialized on the appropriate side of the unstable fixed point, taking care of all off-critical portions of Theorem 1.1.
Criticality. We now turn to the mixing time at the critical point $\beta=\beta_{\mathsf{c}}$. Fix an $\varepsilon$ total-variation distance we are trying to achieve, and in turn take $\gamma$ sufficiently large. By Lemma 2.1, the stationary distribution $\mu$ is a $\xi,1-\xi$ mixture of $\mu^{\mathsf{dis}}$ and $\mu^{\mathsf{ord}}$ for $\xi$ defined in (6). By Lemma 3.22, there is a unique $c_{*}(\xi)$ such that the escape probabilities of Lemma 3.22 are within $o(1)$ of $\xi,1-\xi$. By Lemma 3.23, if $X_{0}\sim G(n,\lambda_{0})$ with $\lambda_{0}=\lambda_{*}+c_{*}n^{-1/2}+o(n^{-1/2})$, then $|\mathbb{P}(\tau_{\gamma}^{X}<\tau_{-\gamma}^{X})-\xi|\leq\varepsilon+o(1)$. Moreover, by Lemma 3.17, the minimum of these two exit times is $O(1)$. Finally, by Lemma 3.14, with probability $1-o(1)$, at exit, the configuration satisfies the necessary conditions to apply Lemma 3.1 and quasi-equilibrate to the phase-restricted measure on the side the dynamics exits with probability $1-O(\gamma^{-2})$. Combined, these imply there exists $T=O(\log n)$ such that if $X_{0}\sim G(n,\lambda_{0})$, then
which will be less than $2\varepsilon$ for $n$ large and $\gamma$ large.
@@ 3.4 Lower bound on the mixing time with different choices of $\lambda_{0}$
We provide the details for item (1), the other cases following by similar reasoning. For any initialization parameter $\lambda_{0}>\lambda_{*}-Kn^{-1/2}$ for some $K=O(1)$, by Lemma 3.22 and 3.23, there is a positive probability $c_{K}>0$ that the process $L_{t}$ hits $\theta_{*}+\gamma n^{-1/2}$ before $\theta_{*}-\gamma n^{-1/2}$ (for sufficiently large $\gamma$) in some $t\leq C_{\gamma}$ many steps. By Lemma 3.14, with probability $1-o(1)$, at exit, the configuration satisfies the necessary conditions to apply Lemma 3.1 and quasi-equilibrate to the right to the ordered phase $\mu^{\mathsf{ord}}$. Putting these together, we find that for some $T_{0}=O(\log n)$,
@@
By using the optimal coupling on these, and then the identity coupling of CM chains after time $T_{0}$,
@@
Next, we claim that by the results of [24], a CM dynamics chain initialized from $\mu^{\mathsf{ord}}$ retains total-variation distance $1-o(1)$ for exponentially many steps to $\mu$ when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$. To see this, note from [24, Lemma 4.7 and the proof of Theorem 2] the existence of a bottleneck set $A$ (that the giant component is at least $(\theta_{*}+\varepsilon)n$ and the number of vertices in components of size larger than $M$ is at most $\rho n$) such that it takes $\exp(\Omega(n))$ steps for a CM dynamics initialized from $\mu(\cdot\mid A)$ to leave $A$. The initialization from $\mu(\cdot\mid A)$ can be seen to be within $e^{-\Omega(n)}$ total-variation distance of an initialization from $\mu^{\mathsf{ord}}$ by an application of Lemma 2.1 and Observation 2.5. From this, we deduce that for every $t$,
@@ 3.5.1 Concentration properties of activation steps
The first equality follows by taking expected values; the second equality is analogous. ∎
@@
$\text{Var}\left(R_{2}^{-}(X^{r})\right)$ $=\mathbb{E}\left[\text{Var}\left(R_{2}^{-}(X^{r})\mid X\right)\right]+\text{Var}\left(\mathbb{E}[R_{2}^{-}(X^{r})\mid X]\right)\leq\mathbb{E}[R_{2}^{-}(X)]+\text{Var}(R_{2}^{-}(X)).$ (49)
@@
$\text{Var}(L_{t+1}\mid{X}_{t},\Lambda_{t+1})=\mathbb{E}[\text{Var}(|\mathcal{L}_{1}(G_{t+1})|\mid A_{t+1})\mid{X}_{t},\Lambda_{t+1}]+\text{Var}(\mathbb{E}[|\mathcal{L}_{1}(G_{t+1})|\mid A_{t+1}]\mid{X}_{t},\Lambda_{t+1})\,.$ (52)
@@
Using that $10<\alpha(\cdot)<$ 1, and taking the variance of the above, we get
@@ 3.5.2 Local limit theorem for the number of activated vertices
$\frac{1}{\log m}\sum_{j=1}^{m}\mathbb{P}(V_{j}=0)\mathbb{P}(V_{j}=M)=\frac{1}{\log m}\sum_{j=1}^{m}\left[\big{(}1-\frac{1}{q}\big{)}\frac{1}{q}\cdot\mathbf{1}[L_{j+1}(X)=M]\right]=O\left(\frac{1}{\log m}\right)\sum_{j=1}^{m}\mathbf{1}[L_{j+1}(X)=M].$
@@
Also, the probability that it doesn’t lie in $[\mu-\sigma n^{1/4},\mu+\sigma n^{1/4}]$ is at most $n^{-1/8}$, say, by Chebyshev’s inequality. Using the fact that $\sigma^{2}=\Theta(n)$, this implies that there exists a coupling $\operatorname{{\mathbf{P}}}$ of $(A^{-},Z)$ such that $\operatorname{{\mathbf{P}}}(|A^{-}-Z|>1)=O(n^{-1/8})$ as desired. ∎

### v2

@@ 3.2.4 Analysis of the limiting 1-dimensional process
Now that we have shown the giant component process near the fixed point $\theta_{*}n$ is well-approximated by the 1-dimensional process $Z_{t}$ of (40), we show here that this simplified process’s exit probabilities to the right and left are monotone, and oscillate on the $\sqrt{n}$ scale. In what follows, let $\tau_{\gamma}^{Z}$ be the hitting time of $\gamma\sqrt{n}$ for $Z_{t}$ from (40) and let $\tau_{-\gamma}^{Z}$ be the hitting time of $\gamma\sqrt{n}$.
@@
For every $p\in(0,1)$, there is a unique $c_{*}\in\mathbb{R}$ such that if $p_{0}=\lambda_{*}+c_{*}n^{-1/2}+o(n^{-1/2})$ and $Z_{0}\sim\mathcal{N}((p_{0}-\lambda_{*})\alpha^{\prime}(\lambda_{*})n,\sigma^{2}(\lambda_{*})n)$ then
@@
We begin by showing that the Markov chain of (40) is monotone in the initialization. Namely, we wish to show that for two initializations $Z_{0}\leq Z_{0}^{\prime}$, the law of $Z_{t}$ is stochastically below that of $Z_{t}^{\prime}$ for all $t$. Suppose that $Z_{t}\leq Z_{t}^{\prime}$ and consider $Z_{t+1}$ and $Z_{t+1}^{\prime}$ generated per (40) using the same Bernoulli random variable $\varepsilon_{t+1}$ and the same pair of normal random variables. If $\varepsilon_{t+1}=0$ then $Z_{t+1}=Z_{t}\leq Z_{t}^{\prime}=Z_{t+1}^{\prime}$ and the monotonicity is preserved. If $\varepsilon_{t+1}=1$, then under this coupling
@@
This will be positive because $Z_{t}^{\prime}-Z_{t}\geq 0$ and $f^{\prime}(\theta_{*})>0$. The monotonicity in the initialization then carries over to monotonicity in $p_{0}$ because $\mathcal{N}(\mu,\sigma^{2})\succeq\mathcal{N}(\mu^{\prime},\sigma^{2})$ if $\mu\geq\mu^{\prime}$.
To see the other consequences, notice first that the process $(\bar{Z}_{t})_{t\geq 0}=(Z_{t}n^{-1/2})_{t\geq 0}$ is $n$-independent (as all of $h_{1},h_{2},\sigma_{t},f^{\prime},\beta_{\mathsf{u}}$ are $n$-independent). This implies that $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ are $n$-independent from $n$-independent initializations, which will be the case if $p_{0}=\lambda_{*}+c_{*}n^{-1/2}$. As $Z_{0}\to\gamma\sqrt{n}$, the probability $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ is easily checked to go to $1$, and as $Z_{0}\to-\gamma\sqrt{n}$, it goes to zero. This implies existence of a unique $c_{*}(\gamma)$ such that from the initialization $Z_{0}$, one has $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})=p+o(1)$, so long as we show continuity of $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ in $c_{*}$. To see that continuity, we show that the total-variation distance between initializations with $p_{0}$ and $p_{0}^{\delta}=p_{0}+\delta n^{-1/2}$ goes to zero as $\delta\downarrow 0$, as there would then be a coupling of the initializations (and therefore the future of the chains including the indicators of which direction it exits) that succeeds with probability going to $1$ as $\delta\downarrow 0$. If $Z_{0}^{\delta}$ has the Gaussian initialization with $p_{0}$ replaced by $p_{0}^{\delta}$, then a total-variation bound between 1-dimensional Gaussians with the same variance gives
$\|\mathbb{P}(Z_{0}\in\cdot)-\mathbb{P}(Z_{0}^{\delta}\in\cdot)\|_{\textsc{tv}}\leq\frac{\delta n^{-1/2}\alpha^{\prime}(\lambda_{*})n}{2\sigma(\lambda_{*})n^{1/2}}=\delta\frac{\alpha^{\prime}(\lambda_{*})}{\sigma(\lambda_{*})}\,$,
which clearly is going to 0 as $\delta\downarrow 0$ as desired.
Finally, to see that $c_{*}$ is $\gamma$-independent up to the $o_{\gamma}(1)$ error, similarly observe that
@@
Suppose $X_{0}\sim G(n,p_{0}/n)$ for $p_{0}=\lambda_{*}+O(n^{-1/2})$ and let $(Z_{t})_{t}$ be the process of Definition 3.18 initialized from $Z_{0}\sim\mathcal{N}((p_{0}-\lambda_{*})\alpha^{\prime}(\lambda_{*})n,\sigma^{2}(\lambda_{*})n)$. Then the hitting probabilities $\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})$ and $\mathbb{P}(\tau_{\gamma}^{X}<\tau_{-\gamma}^{X})$ are within $o_{\gamma,n}(1)$ of one another.
@@
The same holds for hitting $\gamma\sqrt{n}$, as well as for the converse implications, i.e., that if $X$ has hit one side or the other then so has $Z$ or will in the next step. Putting these together, we deduce that for every $\varepsilon$, there exists $\gamma$ sufficiently large such that $|\mathbb{P}(\tau_{\gamma}^{Z}<\tau_{-\gamma}^{Z})-\mathbb{P}(\tau_{\gamma}^{X}<\tau_{-\gamma}^{X})|\leq\varepsilon+o(1)$. ∎
@@ 3.3 Lower bound on the mixing time with different choices of $\lambda_{0}$
We provide the details for item (1), the other cases following by similar reasoning. For any initialization parameter $\lambda_{0}>\lambda_{*}-Kn^{-1/2}$ for some $K=O(1)$, by Lemma 3.22 and 3.23, there is a positive probability $c_{K}>0$ that the process $L_{t}$ hits $\theta_{*}+\gamma n^{-1/2}$ before $\theta_{*}-\gamma n^{-1/2}$ (for sufficiently large $\gamma$) in some $t\leq C_{\gamma}$ many steps. By Lemma 3.14, with probability 1o(1)1-o(1), at exit, the configuration satisfies the necessary conditions to apply Lemma 3.1 and quasi-equilibrate to the right to the ordered phase $\mu^{\mathsf{ord}}$. Putting these together, we find that for some $T_{0}=O(\log n)$,
@@
By using the optimal coupling on these, and then the identity coupling of CM chains after time $T_{0}$, for $t\geq 0$ we have
@@
Next, we claim that by the results of [25], a CM dynamics chain initialized from $\mu^{\mathsf{ord}}$ retains total-variation distance 1o(1)1-o(1) for exponentially many steps to $\mu$ when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$. To see this, note from [25, Lemma 4.7 and the proof of Theorem 2 in the subcritical/critical regime] the existence of a bottleneck set $A$ (that the giant component is at least $(\theta_{*}+\varepsilon)n$ and the number of vertices in non-giant components of size larger than $M$ is at most $\rho n$) such that it takes $\exp(\Omega(n))$ steps for a CM dynamics initialized from $\mu(\cdot\mid A)$ to leave $A$. The initialization from $\mu(\cdot\mid A)$ can be seen to be within $e^{-\Omega(n)}$ total-variation distance of an initialization from $\mu^{\mathsf{ord}}$ by an application of Lemma 2.1 and Observation 2.5. From this, we deduce that for every $t\geq 1$,
@@
3.4 Proof of Theorem 1.1
With the above ingredients at hand, we are in position to complete the proof of Theorem 1.1.
Proof of Theorem 1.1.
We begin by concluding the bounds of mixing when initialized on the appropriate side of the unstable fixed point, taking care of all off-critical portions of Theorem 1.1.
We now turn to the mixing time at the critical point $\beta=\beta_{\mathsf{c}}$. Fix an $\varepsilon$ total-variation distance we are trying to achieve, and in turn take $\gamma$ sufficiently large. By Lemma 2.1, the stationary distribution $\mu$ is a $\xi,1-\xi$ mixture of $\mu^{\mathsf{dis}}$ and $\mu^{\mathsf{ord}}$ for $\xi$ defined in (6). By Lemma 3.22, there is a unique $c_{*}(\xi)$ such that the escape probabilities of Lemma 3.22 are within $o(1)$ of $\xi,1-\xi$. By Lemma 3.23, if $X_{0}\sim G(n,\lambda_{0})$ with $\lambda_{0}=\lambda_{*}+c_{*}n^{-1/2}+o(n^{-1/2})$, then $|\mathbb{P}(\tau_{\gamma}^{X}<\tau_{-\gamma}^{X})-\xi|\leq\varepsilon+o(1)$. Moreover, by Lemma 3.17, the minimum of these two exit times is $O(1)$. Finally, by Lemma 3.14, with probability 1o(1)1-o(1), at exit, the configuration satisfies the necessary conditions to apply Lemma 3.1 and quasi-equilibrate to the phase-restricted measure on the side the dynamics exits with probability $1-O(\gamma^{-2})$. Combined, these imply there exists $T=O(\log n)$ such that if $X_{0}\sim G(n,\lambda_{0})$, then
which will be less than $2\varepsilon$ for $n$ large and $\gamma$ large.
The slow mixing claims are exactly the statement of Theorem 3.24. ∎
We also include a proof of Theorem 1.3 on initializations from the random-cluster Gibbs measure at a different temperature, to justify why it follows from the above arguments.
Proof of Theorem 1.3.
It is evident at this point that the only properties of the initialization used in our proof of Theorem 1.1 were on the size of its giant component in relation to $\theta_{*}n$, its second largest component size, and the sum of squares of its non-giant components. The requisite properties on the latter two quantities are known to hold for samples from $\mu_{\beta_{0}}$ per Lemma 2.1. The requisite property of the giant component size drawn from $\mu_{\beta_{0}}$ is that it should be on the side of $\theta_{*}$ to which it aims to quasi-equilibrate.
In item 1 when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, we let $b_{*}(\beta,q)=\beta_{\mathsf{c}}$. To see this, first note that if $\beta_{0}<\beta_{\mathsf{c}}$, then the size of the largest component in a sample from $\mu_{\beta_{0}}$ is $O(\log n)$ with high probability and from such initialization the CM dynamics will mix in $O(\log n)$ steps. On the other hand, if $\beta_{0}>\beta_{\mathsf{c}}$, then a typical sample from $\mu_{\beta_{0}}$ has the largest component of size close to $\theta_{\mathsf{r}}(\beta_{0})n$, where we recall from Lemma 2.1 and Lemma 2.12 that $\theta_{\mathsf{r}}(b)$ is the typical size of a giant in a sample from $\mu_{b}$ as well as the second root of $f$. By monotonicity, $\theta_{\mathsf{r}}(\beta_{0})>\theta_{\mathsf{r}}(\beta)>\theta_{*}(\beta)$, and hence the CM dynamics initialized from $\mu_{\beta_{0}}$ has exponential slow mixing. In item 2, if $\beta_{0}>\beta_{\mathsf{c}}$ then $\theta_{\mathsf{r}}(\beta_{0})>\theta_{\mathsf{r}}(\beta_{\mathsf{c}})>\theta_{*}(\beta_{c})$ by monotonicity; also, if $\beta_{0}<\beta_{c}$ then the CM dynamics will start with a configuration from $\mu_{\beta_{0}}$, which has no large components. In either case, the size of largest component of the starting configuration is not close to $\theta_{*}(\beta)$, so there is no fast mixing. Finally, in item 3 when $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$, we define $b_{*}(\beta,q)=\inf\{b:\theta_{\mathsf{r}}(b)>\theta_{*}(\beta)\}$ such that $O(\log n)$ mixing occurs if and only if $\beta_{0}>b_{*}(\beta,q)$. ∎
@@ 3.5.1 Concentration properties of activation steps
The first claim of the observation then follows by taking expected values and the second claim is analogous. ∎
@@
$\text{Var}\left(R_{2}^{-}(X^{r})\right)$ $=\mathbb{E}\left[\text{Var}\left(R_{2}^{-}(X^{r})\mid X\right)\right]+\text{Var}\left(\mathbb{E}[R_{2}^{-}(X^{r})\mid X]\right)$
$\leq\mathbb{E}[R_{2}^{-}(X)]+\text{Var}(R_{2}^{-}(X)).$ (49)
@@
$\text{Var}(L_{t+1}\mid$ ${X}_{t},\Lambda_{t+1})$
$=\mathbb{E}[\text{Var}(|\mathcal{L}_{1}(G_{t+1})|\mid A_{t+1})\mid{X}_{t},\Lambda_{t+1}]+\text{Var}(\mathbb{E}[|\mathcal{L}_{1}(G_{t+1})|\mid A_{t+1}]\mid{X}_{t},\Lambda_{t+1})\,.$ (52)
@@
Using that $0<\alpha(\cdot)<1$, and taking the variance of the above, we get
@@ 3.5.2 Local limit theorem for the number of activated vertices
$\frac{1}{\log m}\sum_{j=1}^{m}\mathbb{P}(V_{j}=0)\mathbb{P}(V_{j}=M)$ $=\frac{1}{\log m}\sum_{j=1}^{m}\left[\big{(}1-\frac{1}{q}\big{)}\frac{1}{q}\cdot\mathbf{1}[L_{j+1}(X)=M]\right]$
$=O\left(\frac{1}{\log m}\right)\sum_{j=1}^{m}\mathbf{1}[L_{j+1}(X)=M].$
@@
Also, the probability that it doesn’t lie in $\mu-\sigma n^{1/4},\mu+\sigma n^{1/4}]$ is at most $n^{-1/8}$, say, by Chebyshev’s inequality. Using the fact that $\sigma^{2}=\Theta(n)$, this implies that there exists a coupling $\operatorname{{\mathbf{P}}}$ of $(A^{-},Z)$ such that $\operatorname{{\mathbf{P}}}(|A^{-}-Z|>1)=O(n^{-1/8})$ as desired. ∎